A soft drink machine is regulated so that it discharges an average of 8 ounces per cup. If the amount of drink per cup is normally distributed with a standard deviation of 0.5 ounce, find the probability a cup contains:
7. More than 8 ounces ____________
8. Between 7.8 and 8.4 ounces ____________
the formulas I seem to be using are not correct in Excel! Please help thank you

Answer by stanbon(75887)

A soft drink machine is regulated so that it discharges an average of 8 ounces per cup. If the amount of drink per cup is normally distributed with a standard deviation of 0.5 ounce, find the probability a cup contains:
7. More than 8 ounces = 0.5
-----------------------------------
8. Between 7.8 and 8.4 ounces
z(7.8) = (7.8-8)/0.5 = -0.2/0.5 = -0.4
z(8.4) = (8.4-8)/0.5 = 0.4/0.5 = 0.8
---
P(7.8 < x < 8.4) = P(-0.4< z <0.8) = normalcdf(-0.4,0.8) = 0.4436
----------------------
Cheers,
Stan H.
===================